# DISTA-Net: Dynamic Closely-Spaced Infrared Small Target Unmixing

### v1

@@ Abstract
Resolving closely-spaced small targets in dense clusters presents a significant challenge in infrared imaging, as the overlapping signals hinder precise determination of their quantity, sub-pixel positions, and radiation intensities. While deep learning has advanced the field of infrared small target detection, its application to closely-spaced infrared small targets has not yet been explored. This gap exists primarily due to the complexity of separating superimposed characteristics and the lack of an open-source infrastructure. In this work, we propose the Dynamic Iterative Shrinkage Thresholding Network (DISTA-Net), which reconceptualizes traditional sparse reconstruction within a dynamic framework. DISTA-Net adaptively generates convolution weights and thresholding parameters to tailor the reconstruction process in real time. To the best of our knowledge, DISTA-Net is the first deep learning model designed specifically for the unmixing of closely-spaced infrared small targets, achieving superior sub-pixel detection accuracy. Moreover, we have established the first open-source ecosystem to foster further research in this field. This ecosystem comprises three key components: (1) CSIST-100K, a publicly available benchmark dataset; (2) CSO-mAP, a custom evaluation metric for sub-pixel detection; and (3) GrokCSO, an open-source toolkit featuring DISTA-Net and other models. Our code and dataset are available at https://github.com/GrokCV/GrokCSO.
@@ 1 Introduction
Infrared imaging plays a pivotal role in various long-distance detection and surveillance tasks Sobrino et al. (2016), due to its exceptional sensitivity to thermal radiation and independence from illumination conditions. However, the radiation intensity captured from remote targets is inherently weak due to their long-range distance from the imaging system Kou et al. (2023). This challenge is exacerbated when targets appear in spatially close dense clusters, as the closely-spaced objects (CSO) Pruett et al. (2023), resulting in overlapping blob-like spots.
As shown in Fig. 1, such overlap makes it impossible to resolve the targets independently via human vision, thus obscuring the perception of target count, precise locations, and radiation intensities Zhao et al. (2022), presenting a substantial impediment for Infrared Search and Tracking (IRST) systems in their subsequent phases of detection, tracking, and identification. Therefore, exploring effective techniques for unmixing and reconstruction of such closely-spaced infrared small targets (CSIST), to accurately discern their exact locations and radiant intensities, holds great significance.
@@
Despite the critical importance of closely-spaced infrared small target unmixing in various applications, research addressing this specific task remains exceedingly scarce. These approaches typically formulate the problem as a parameter estimation task and employ optimization algorithms to solve it Meyer and Williams (2021). Target sparsity on the imaging plane was leveraged to devise a discretized sampling-based sparse reconstruction method Zhang et al. (2013), utilizing an over-complete dictionary and solving a second-order cone programming problem under the $\ell_{1}$ norm regularization.
@@
However, the performance of these optimization-based models is highly dependent on meticulous hyperparameter tuning Chen et al. (2021), which poses significant challenges in real-world scenarios. Variations in target quantity or location further complicate the selection of optimal hyperparameters, limiting the generalizability and practicality of these methods Wan et al. (2018). Consequently, there is a pressing need for the development of unmixing algorithms that exhibit greater robustness to hyperparameter variations and can be effectively applied in diverse real-world settings.
@@
While deep learning has revolutionized image super-resolution Li et al. (2023), its application to closely-spaced infrared small target unmixing remains unexplored, primarily due to fundamentally different task objectives and ecosystem limitations. Unlike generic super-resolution that enhances clarity through high-frequency detail restoration, infrared small target unmixing requires precise estimation of overlapping targets’ count, locations, and radiation intensities - essentially mapping images to specific target attributes rather than conventional pixel-space super-resolution. This challenge is further compounded by the absence of standardized benchmark datasets, task-specific evaluation metrics, and open-source implementations, creating significant barriers to developing and comparing deep learning approaches in this specialized domain.
@@
To address the aforementioned challenges, in this paper, we propose a novel deep unfolding network for closely-spaced infrared small target unmixing, termed as the Dynamic Iterative Shrinkage Thresholding Network (DISTA-Net). We reformulate the traditional sparse reconstruction approach into a dynamic deep learning framework, which adaptively generates convolution weights and thresholding parameters to tailor the reconstruction process in real time. Distinct from prior methods, the parameters associated with the proximal mapping, including nonlinear transforms and shrinkage thresholds, are dynamically adapted to the input data, rather than being hand-crafted or fixed after training. To the best of our knowledge, this represents the first deep learning-based effort towards unmixing closely-spaced infrared small targets.
@@
Furthermore, we have established a comprehensive open-source ecosystem to facilitate research in this domain. This ecosystem includes CSIST-100K (Closely-Spaced Infrared Small Target), an open benchmark dataset comprising 100,000 pairs of low-resolution images and exact annotations of location and radiation intensities. We also introduce CSO-mAP, a custom evaluation metric inspired by the mean average precision (mAP) from object detection, calibrated to evaluate the quantity, spatial positioning, and radiation intensity of the unmixed infrared small targets. Accompanying these resources is GrokCSO, an open-source PyTorch-based toolkit encapsulating our DISTA-Net alongside other state-of-the-art models, empowering researchers to effortlessly leverage the CSIST-100K dataset.
In summary, our contributions can be categorized into Four main aspects:
  1. 1.
    We reformulate the unmixing of closely-spaced infrared small targets as an interpretable deep unfolding problem. To our knowledge, this is the first deep learning based attempt for this task.
  2. 2.
    Our proposed DISTA-Net is a dynamic deep unfolding network, which adaptively generates convolution weights and thresholding parameters to tailor the reconstruction process conditioned on the input data.
  3. 3.
    We establish the first open-source ecosystem for this task, including the CSIST-100K dataset, the CSO-mAP metric, and the GrokCSO toolkit.
  4. 4.
    We provide a comprehensive analysis of our approach, validating the importance of dynamic deep unfolding and the effectiveness of the DISTA-Net in addressing the CSIST unmixing task.
2 Relate Work
@@ 2.1 Infrared Small Target Detection
Driven by a range of open-source datasets Dai et al. (2021a, 2023b); Li et al. (2024b), infrared small target detection has garnered significant research attention in recent years. Current research mainly focuses on developing multi-scale feature fusion models to counteract the scarcity of intrinsic target features Xu et al. (2023). Dai et al. introduced an asymmetric contextual modulation module that bridges high-level semantics with low-level details using a bottom-up pathway via point-wise channel attention Dai et al. (2021b). Wang et al. merged reinforcement learning with pyramid feature fusion and proposed a global context boundary attention module to mitigate localized bright noise Wang et al. (2023). Tong et al. adopted an encoder–decoder structure, enhancing feature extraction with an atrous spatial pyramid pooling and a dual-attention module, and using multiscale labels to focus on target edges and internal features Tong et al. (2023).
Our work focuses on infrared small targets but differs from previous studies in two main ways. First, while infrared small target detection is a precursor to our study, our primary focus is on the unmixing of closely-spaced infrared small targets. Detecting potentially overlapped targets is essential for subsequent tasks like sub-pixel localization and radiation intensity prediction. Second, our task extends beyond simple detection by precisely locating targets at the sub-pixel level and estimating their radiative intensities, offering a more detailed understanding of target characteristics compared to the previous binary detection approach.
@@ 2.2 Deep Unfolding
Deep unfolding, as delineated in Monga et al. (2021), originated in 2010 with Learned Iterative Shrinkage-Thresholding Algorithm (LISTA) Gregor and LeCun (2010), which reinterprets ISTA Daubechies et al. (2004) as a fully connected feed-forward neural network. This approach generalizes effectively to new samples, achieving ISTA-like accuracy with fewer iterations. Subsequent works, such as ADMM-Net yang et al. (2016), have unfolded the steps of the Alternating Direction Method of Multipliers (ADMM) into a deep learning framework, thereby improving the accuracy and efficiency of MRI reconstruction through a compressive sensing model optimized via end-to-end discriminative training. Likewise, ISTA-Net Zhang and Ghanem (2018) has adopted an end-to-end learning approach for proximal mapping, enhancing the performance of compressive sensing for natural image reconstruction.
Motivated by such advances, deep unfolding has found applicability in a variety of computer vision tasks. Notably, Li et al. transformed a generalized gradient-domain total variation algorithm into a deep interpretable network for blind image deblurring, delivering superior performance with learned parameters Li et al. (2020). For image super-resolution, Guo et al. incorporated trainable convolutional layers into the Discrete Cosine Transform framework, effectively mitigating artifacts and enabling learning from limited data Guo et al. (2019). Solomon et al. unfolded robust principal component analysis into a deep network, improving the distinction between microbubble and tissue signals in ultrasound imaging Solomon et al. (2020).
Unlike previous methods with static parameters Zhang and Ghanem (2018); You et al. (2021), our DISTA-Net dynamically adapts proximal mapping weights based on the input, enabling an adaptive reconstruction process that caters to varying scenarios.
@@ 3 CSIST Benchmark, Metric, and Toolkit
Figure 2: CSIST Visualization: The top row shows 1 to 5 overlapping targets, and the following rows display unmixing results for sub-pixel division factors of 3×, 5×, and 7×.
CSIST-100K Dataset. When a single point source is imaged through a lens, an Airy pattern emerges at the focal point, characterized by a central bright spot comprising 84% of the total energy, surrounded by multiple concentric diffraction rings. This diffraction effect is typically approximated by the point spread function (PSF), whose standard deviation, dependent on the sensor’s focal ratio ($f$-number) and detection band, quantifies the energy spread. In multi-target imaging, the intensity captured by each pixel is the cumulative response of the overlaid point sources. Particularly for remote objects, each is considered a point source, with the radius of the resultant Airy spot dictated by $1.22\lambda/D$, where $\lambda$ is the wavelength and $D$ the lens diameter. This radius is equivalent to $1.9\sigma$ of a Gaussian PSF, marking the sensor’s physical resolution limit, as per the Rayleigh unit.
In our study, we set $\sigma_{\mathrm{PSF}}$ at 0.5 pixels. Our simulations include 1 to 5 targets per image, each characterized by sub-pixel coordinates and radiation intensity. Targets with radiation intensities between 220 and 250 units are randomly positioned within a single pixel in an $11\times 1$ 11 pixel grid, while maintaining separation of at least 0.52 Rayleigh units. We generated the CSIST-100K dataset with 100,000 samples, using 80,000 for training and dividing the remaining 20,000 equally between validation and testing. As shown in Fig. 2, closely-spaced targets undergo diffusion within the pixel matrix, with energy primarily concentrated in a $3\times 3$ pixel area. This creates significant overlap that makes it extremely difficult to determine both the number of distinct targets and their precise spatial coordinates.
CSO-mAP Metric.
Traditional bounding box metrics fails when targets lie within $1.9\sigma$ (due to Airy spot interference), Our CSO-mAP with strict sub-pixel position/intensity matching for CSIST evaluation, redefining TP/FP as follows:
$\mathbbm{1}_{k}\left(\hat{t}_{j},t_{i}\right)=\left\{\begin{array}[]{lc}1,&\text{ if }d\left(\hat{t}_{j},t_{i}\right)<\delta_{k}\\ 0,&\text{ otherwise }\end{array}\right.$ (1)
Here, $\hat{t}_{j}$ is the $j$-th predicted target, $t_{i}$ is the $i$-th true target, and $\delta_{k}\in\left\{0.05,0.1,0.15,0.2,0.25\right\}$ is a series of distance thresholds, with $k=1,2,3,4,5$, used to control the desired localization accuracy. Precision-recall (PR) curves are generated through intensity-prioritized matching, with Average Precision (AP) computed at each $\delta_{k}$. The final CSO-mAP metric is derived by averaging AP values across all thresholds, explicitly quantifying performance under varying spatial resolution demands, see Supplementary for details.
GrokCSO Toolkit. To address the lack of specialized tools in this domain, we introduce GrokCSO, an open-source toolkit for enhancing the reconstruction of closely-spaced infrared small targets. Built on PyTorch, GrokCSO provides pre-trained models, reproducibility scripts, and specialized evaluation metrics tailored for CSO challenges. The detailed architecture and features of this toolkit are provided in Supplementary.
@@ 4.1 Imaging and Unmixing Framework
CSIST Imaging Model.
Given the significant distance between targets and the infrared detector, targets can be approximated as point sources. The optical system’s diffraction spreads the energy across adjacent pixels, described by a two-dimensional Gaussian point spread function (PSF) Liaudat et al. (2021):
@@
where $\sigma_{\mathrm{PSF}}$ represents the diffusion variance, and $(x_{i},y_{i})$ denote the target’s coordinates. For $N$ closely-spaced targets, each target is represented as $t_{i}=(x_{i},y_{i},g_{i},\sigma_{\mathrm{PSF}})$, where $i\in{1,..,N}$, with $(x_{i},y_{i})$ denoting sub-pixel coordinates and $g_{i}$ the radiation intensities. The collections of center points and intensities are denoted as $\mathcal{P}=\left\{p_{i}|i\in\{1,..,N\}\right\}$ and $\mathcal{G}=\left\{g_{i}|i\in\{1,..,N\}\right\}$. On an infrared focal plane of $U\times V$ pixels, each pixel integrates the PSF within its boundaries:
$g_{i,j}\left(x_{t},y_{t}\right)=\int_{x_{i,j}-1/2D}^{x_{i,j}+1/2D}\int_{y_{i,j}-1/2D}^{y_{i,j}+1/2D}p(x,y)\mathrm{d}x\mathrm{\leavevmode\nobreak\ d}y,$ (3)
where $(x_{i,j},y_{i,j})$ is the pixel’s center, $i=0,…,U,j=0,…,V-1$, with $D$ being the pixel width. The focal plane measurement model is vectorized as:
@@
where $\mathbf{G}(\mathbf{x},\mathbf{y})$ is the steering vector matrix, $\mathbf{s}$ represents target intensities, and $\mathbf{n}$ denotes Gaussian white noise.
@@
As shown in Fig. 3, each pixel is divided into an $n\times n$ grid, resulting in $L=UVn^{2}$ sub-pixels. With sufficient grid resolution, each sub-pixel contains at most one target, with maximum position deviation of $\sqrt{2}D/n$. The measurement model can be extended to $\Omega$:
@@
where $\mathbf{G}(\Omega)$ contains steering vectors from $\Omega$, $\mathbf{w}$ denotes Gaussian white noise, and $\tilde{\mathbf{s}}\in\mathrm{R}^{L}$ is sparse with $L\gg UV$. The CSIST unmixing problem can then be formulated as a sparse reconstruction with $\ell_{1}$ regularization:
@@
Optimization Solution. The ISTA Daubechies et al. (2004) solves the sparse reconstruction problem through two alternating steps:
@@
$\tilde{\mathbf{s}}^{(k)}$ $=\underset{\tilde{\mathbf{s}}}{\arg\min}\frac{1}{2}\left\|\tilde{\mathbf{s}}-\mathbf{r}^{(k)}\right\|_{2}^{2}+\lambda\|\boldsymbol{\Psi}\tilde{\mathbf{s}}\|_{1}.$ (8)
where $\boldsymbol{\Psi}$ denotes the transform matrix, $k$ the iteration index, and $\rho$ the step size. The second step represents a proximal mapping:
$\operatorname{prox}_{\lambda\phi}(\mathbf{r})=\arg\min_{\tilde{\mathbf{s}}}\frac{1}{2}\|\tilde{\mathbf{s}}-\mathbf{r}\|_{2}^{2}+\lambda\phi(\tilde{\mathbf{s}}).$ (9)
@@
Figure 4: Architecture of the proposed DISTA-Net. The overall framework consists of multiple cascaded stages. Each stage contains three main components: a dual-branch dynamic transform module ($\mathcal{F}_{d}^{(k)}$) for feature extraction, a dynamic threshold module ($\Theta_{d}^{(k)}$) for feature refinement, and an inverse transform module ($\tilde{\mathcal{F}_{d}}^{(k)}$) for reconstruction.

### v2

@@ Abstract
Resolving closely-spaced small targets in dense clusters presents a significant challenge in infrared imaging, as the overlapping signals hinder precise determination of their quantity, sub-pixel positions, and radiation intensities. While deep learning has advanced the field of infrared small target detection, its application to closely-spaced infrared small targets has not yet been explored. This gap exists primarily due to the complexity of separating superimposed characteristics and the lack of an open-source infrastructure. In this work, we propose the Dynamic Iterative Shrinkage Thresholding Network (DISTA-Net), which reconceptualizes traditional sparse reconstruction within a dynamic framework. DISTA-Net adaptively generates convolution weights and thresholding parameters to tailor the reconstruction process in real time. To the best of our knowledge, DISTA-Net is the first deep learning model designed specifically for the unmixing of closely-spaced infrared small targets, achieving superior sub-pixel detection accuracy. Moreover, we have established the first open-source ecosystem to foster further research in this field. This ecosystem comprises three key components: (1) CSIST-100K, a publicly available benchmark dataset; (2) CSO-mAP, a custom evaluation metric for sub-pixel detection; and (3) GrokCSO, an open-source toolkit featuring DISTA-Net and other state-of-the-art models, available at https://github.com/GrokCV/GrokCSO.
@@ 1 Introduction
Infrared imaging plays a pivotal role in various long-distance detection and surveillance tasks [26], due to its exceptional sensitivity to thermal radiation and independence from illumination conditions. However, the radiation intensity captured from remote targets is inherently weak due to their long-range distance from the imaging system [11]. This challenge is exacerbated when targets appear in spatially close dense clusters, as the closely-spaced objects (CSO) [24], resulting in overlapping blob-like spots.
As shown in Fig. 1, such overlap makes it impossible to resolve the targets independently via human vision, thus obscuring the perception of target count, precise locations, and radiation intensities [46], presenting a substantial impediment for Infrared Search and Tracking (IRST) systems in their subsequent phases of detection, tracking, and identification. Therefore, exploring effective techniques for unmixing and reconstruction of such closely-spaced infrared small targets (CSIST), to accurately discern their exact locations and radiant intensities, holds great significance.
@@
Despite the critical importance of CSIST unmixing in various applications, research addressing this specific task remains exceedingly scarce. These approaches typically formulate the problem as a parameter estimation task and employ optimization algorithms to solve it [21]. Target sparsity on the imaging plane was leveraged to devise a discretized sampling-based sparse reconstruction method [39], utilizing an over-complete dictionary and solving a second-order cone programming problem under the $\ell_{1}$ norm regularization. However, the performance of these optimization-based models is highly dependent on meticulous hyperparameter tuning [1], which poses significant challenges in real-world scenarios. Variations in target quantity or location further complicate the selection of optimal hyperparameters, limiting the generalizability and practicality of these methods [29]. Consequently, there is a pressing need for the development of unmixing algorithms that exhibit greater robustness to hyperparameter variations and can be effectively applied in diverse real-world settings.
@@
While deep learning has revolutionized image super-resolution [13], its application to CSIST unmixing remains unexplored, primarily due to fundamentally different task objectives and ecosystem limitations. Unlike generic super-resolution that enhances clarity through high-frequency detail restoration [33], infrared small target unmixing requires precise estimation of overlapping targets’ counts, locations, and radiation intensities - essentially mapping images to specific target attributes rather than conventional pixel-space super-resolution. This challenge is further compounded by the absence of standardized benchmark datasets, task-specific evaluation metrics, and open-source implementations, creating significant barriers to developing and comparing deep learning approaches in this specialized domain.
@@
To address the aforementioned challenges, in this paper, we propose a novel deep unfolding network for CSIST unmixing, termed as the Dynamic Iterative Shrinkage Thresholding Network (DISTA-Net). We reformulate the traditional sparse reconstruction approach into a dynamic deep learning framework, which adaptively generates convolution weights and thresholding parameters to tailor the reconstruction process in real time. Distinct from prior methods, the parameters associated with the proximal mapping (nonlinear transforms and shrinkage thresholds) are dynamically adapted to the input data, rather than being hand-crafted or fixed after training. To the best of our knowledge, this represents the first deep learning-based effort towards CSIST unmixing.
@@
Furthermore, we establish a comprehensive open-source ecosystem to facilitate research in this domain, including CSIST-100K, an open benchmark dataset comprising 100,000 pairs of CSIST images and exact annotations of location and radiation intensities; CSO-mAP, a custom evaluation metric inspired by the mean average precision (mAP) from object detection, calibrated to evaluate the quantity, spatial positioning, and radiation intensity of the unmixed infrared targets; and GrokCSO, an open-source PyTorch-based toolkit encapsulating our DISTA-Net alongside other state-of-the-art models, empowering researchers to effortlessly leverage the CSIST-100K dataset.
Our contributions can be categorized into Four aspects: 1. We reformulate CSIST unmixing as an interpretable deep unfolding problem. To our knowledge, this is the first deep learning based attempt for this task. 2. Our proposed DISTA-Net is a dynamic deep unfolding network, which adaptively generates convolution weights and thresholding parameters to tailor the reconstruction process conditioned on the input data. 3. We establish the first open-source ecosystem for this task, including the CSIST-100K dataset, the CSO-mAP metric, and the GrokCSO toolkit. 4. We provide a comprehensive analysis of our approach, validating the importance of dynamic deep unfolding and the effectiveness of the DISTA-Net in addressing the CSIST unmixing task.
2 Related Work
@@ 2.1 Infrared Small Target Detection
Driven by a range of open-source datasets [4, 6, 15], infrared small target detection has garnered significant research attention in recent years. Current research mainly focuses on developing multi-scale feature fusion models to counteract the scarcity of intrinsic target features [35]. Dai et al. introduced an asymmetric contextual modulation module that bridges high-level semantics with low-level details using a bottom-up pathway via point-wise channel attention [5]. Wang et al. merged reinforcement learning with pyramid feature fusion and proposed a global context boundary attention module to mitigate localized bright noise [32]. Cheng et al. proposed a difference-aware attention module with a dual-temporal aggregation module for global feature learning and channel activation, and a difference-attention module for multi-scale detection via local correlations [20]. Tong et al. adopted an encoder–decoder structure, enhancing feature extraction with an atrous spatial pyramid pooling and a dual-attention module, and using multiscale labels to focus on target edges and internal features [28]. Zhang et al. proposed an infrared small target detection framework integrating visual-textual information via CLIP-prompted SAM adaptation with a denoising module, achieving enhanced generalization capability for the infrared domain [41].
Our work focuses on infrared small targets but differs in two key aspects. First, while infrared small target detection precedes our study, we prioritize CSIST unmixing, where detecting overlapped targets is crucial for sub-pixel localization and radiation intensity prediction. Second, our task goes beyond detection by enabling sub-pixel-level localization and radiative intensity estimation, providing finer target characterization than binary detection.
@@ 2.2 Deep Unfolding
Deep unfolding, as delineated in [22], originated in 2010 with the Learned Iterative Shrinkage-Thresholding Algorithm (LISTA) [9], which reinterprets ISTA [7] as a fully connected feed-forward neural network. This approach generalizes effectively to new samples, achieving ISTA-like accuracy with fewer iterations. Subsequent works, such as ADMM-Net [36], have unfolded the steps of the Alternating Direction Method of Multipliers (ADMM) into a deep learning framework, thereby improving the accuracy and efficiency of MRI reconstruction through a compressive sensing model optimized via end-to-end discriminative training. Likewise, ISTA-Net [40] has adopted an end-to-end learning approach for proximal mapping, enhancing the performance of compressive sensing for natural image reconstruction.
Motivated by such advances, deep unfolding has found applicability in a variety of computer vision tasks. Notably, Li et al. transformed a generalized gradient-domain total variation algorithm into a deep interpretable network for blind image deblurring, delivering superior performance with learned parameters [14]. For image super-resolution, Guo et al. incorporated trainable convolutional layers into the Discrete Cosine Transform framework, effectively mitigating artifacts and enabling learning from limited data [10]. Solomon et al. unfolded robust principal component analysis into a deep network, improving the distinction between microbubble and tissue signals in ultrasound imaging [27].
Unlike previous methods with static parameters [40, 37], our DISTA-Net dynamically adapts proximal mapping weights based on the input, enabling an adaptive reconstruction process that caters to varying scenarios.
@@ 3 CSIST Benchmark, Metric, and Toolkit
Figure 2: CSIST Visualization: The top row shows 1 to 5 overlapping targets, and the following rows display unmixing results for sub-pixel division factors of $3\times$, $5\times$, and $7\times$.
CSIST-100K Dataset. In our study, we set $\sigma_{\mathrm{PSF}}$ at 0.5 pixels. Simulations include 1–5 overlapping targets per image, each defined by 2D coordinates and radiation intensity (220–250 units), randomly placed within an $11\times 11$ grid while maintaining $\geq$ 0.52 Rayleigh units separation. We generated the CSIST-100K dataset with 100,000 samples: 80,000 for training, and 20,000 split equally for validation and testing. As shown in Fig. 2, closely spaced targets diffuse, with energy concentrated in a $3\times 3$ pixel area, causing significant overlap that complicates target counting and coordinate determination. (See Supplementary for optical modeling details.)
CSO-mAP Metric. Traditional bounding box metrics fail when target separation falls below the Rayleigh criterion (due to Airy spot interference). Our CSO-mAP with strict sub-pixel position/intensity matching for CSIST evaluation, redefining TP/FP as follows:
[…]
Here, $\delta_{k}\in\left\{0.05,0.1,0.15,0.2,0.25\right\}$ is a series of distance thresholds, with $k=1,2,3,4,5$, used to control the desired localization accuracy. Precision-recall (PR) curves are generated through intensity-prioritized matching, with Average Precision (AP) computed at each $\delta_{k}$. The final CSO-mAP metric is derived by averaging AP values across all thresholds, explicitly quantifying performance under varying spatial resolution demands (details in Supplementary).
GrokCSO Toolkit. To address the lack of specialized tools in this domain, we introduce GrokCSO, an open-source toolkit for CSIST unmixing. Built on PyTorch, GrokCSO provides pre-trained models, reproducibility scripts, and specialized evaluation metrics tailored for CSO challenges. The detailed architecture and features of this toolkit are provided in Supplementary.
@@ 4.1 Imaging and Unmixing Framework
CSIST Imaging Model. Given the significant distance between targets and the infrared detector, targets can be approximated as point sources. The optical system’s diffraction spreads the energy across adjacent pixels, described by a two-dimensional Gaussian point spread function (PSF) [18]:
@@
where $\sigma_{\mathrm{PSF}}^{2}$ is the diffusion variance and $(x_{t},y_{t})$ the target coordinates. On an infrared focal plane of $U\times V$ pixels, each pixel integrates the PSF within its boundaries:
$g_{i,j}\left(x_{t},y_{t}\right)=\int_{x_{i,j}-1/2D}^{x_{i,j}+1/2D}\int_{y_{i,j}-1/2D}^{y_{i,j}+1/2D}p(x,y)\mathrm{d}x\mathrm{~d}y,$ (3)
where $(x_{i,j},y_{i,j})$ is the pixel’s center and $D$ the pixel width. The focal plane measurement model is vectorized as:
@@
where $\mathbf{G}(\mathbf{x},\mathbf{y})$ is the steering matrix, $\mathbf{s}$ represents target intensities, and $\mathbf{n}$ denotes Gaussian white noise.
@@
As shown in Fig. 3, each pixel is divided into an $n\times n$ grid, resulting in $L=UVn^{2}$ sub-pixels. With sufficient grid resolution, each sub-pixel contains at most one target, with maximum position deviation of $\sqrt{2}D/2n$. The measurement model can be extended to $\Omega$:
@@
where $\mathbf{G}(\Omega)$ contains steering vectors from $\Omega$, and $\tilde{\mathbf{s}}\in\mathrm{R}^{L}$ is sparse with $L\gg UV$. The CSIST unmixing problem can then be formulated as a sparse reconstruction with $\ell_{1}$ regularization:
@@
Optimization Solution. The ISTA [7] solves the sparse reconstruction problem through two alternating steps:
@@
$\tilde{\mathbf{s}}^{(k)}$ $=\underset{\tilde{\mathbf{s}}}{\arg\min}\frac{1}{2}\left\|\tilde{\mathbf{s}}-\mathbf{r}^{(k)}\right\|_{2}^{2}+\lambda\|\bm{\Psi}\tilde{\mathbf{s}}\|_{1},$ (8)
where $\bm{\Psi}$ denotes the transform matrix, $k$ the iteration index, and $\rho$ the step size. The second step represents a proximal mapping:
[…]
@@
Figure 4: Architecture of the proposed DISTA-Net. The overall framework consists of multiple cascaded stages. Each stage contains three main components: a dual-branch dynamic transform module ($\mathcal{F}_{d}^{(k)}$) for feature extraction, a dynamic threshold module ($\Theta_{d}^{(k)}$) for feature refinement, and an inverse transform module ($\tilde{\mathcal{F}}^{(k)}$) for reconstruction.